OpenStudy (anonymous):

Trigonometry Help, just a bit confused with one part @ganeshie8

OpenStudy (anonymous):

Ill post my work on it so far and you can see if ive done it right but im as far as i can go without getting stuck

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

my first step^

ganeshie8 (ganeshie8):

hold up

OpenStudy (anonymous):

kk

ganeshie8 (ganeshie8):

\[\large \sec^2x - 2\tan^2x = -2\]

ganeshie8 (ganeshie8):

\[\large 1+ \tan^2 x - 2\tan^2x = -2\]

ganeshie8 (ganeshie8):

\[\large 1-\tan^2x = -2\]

ganeshie8 (ganeshie8):

\[\large \tan^2x = 3\]

ganeshie8 (ganeshie8):

\[\large \tan x = \pm \sqrt{3}\]

ganeshie8 (ganeshie8):

this way, it looks bit simple right ?

OpenStudy (anonymous):

oh wow yes i see exactly what you did. cant believe i didnt see that

OpenStudy (anonymous):

if i woulda done it the long way, i would still come out wiht the same answer though right?

ganeshie8 (ganeshie8):

yes ! all roads lead to the same destination

OpenStudy (anonymous):

awesome! wow that problem was so so so much easier your way ahha

OpenStudy (ikram002p):

all roads lea to Roma hehe

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

alright so i get pi/3, 5pi/3, 5pi/3 do i need to put 5pi/3 twice or just put it once?

ganeshie8 (ganeshie8):

lol okay finding solutions for \(\large \tan x = \pm \sqrt{3}\) can be tricky as the period is NOT same as sin or cos... keepin mind the period of tan is pi, NOT 2pi.

OpenStudy (anonymous):

oh goodness so i have to do it a diff way? i cant just do inverse tan of sqrt 3?

ganeshie8 (ganeshie8):

yes, instead of adding 2pi, you need to add pi

OpenStudy (anonymous):

ooh i use the tan x = sinx/cosx?

OpenStudy (anonymous):

so it would be (sin sqrt3)/(cos sqrt3)?

ganeshie8 (ganeshie8):

\(\large \tan x = \sqrt{3}\) : since \(\pi/3\) is one solution, pi/3 + pi = \(4\pi/3\) is also a solution

OpenStudy (anonymous):

ok so i was totally wrong

OpenStudy (anonymous):

how do we know Pi/3 is a solution exactly?

ganeshie8 (ganeshie8):

short answer : we knw \(\pi/3\) is a solution cuz \(\large \tan(\pi/3) = \sqrt{3}\)

OpenStudy (anonymous):

hmm i think i udnerstand that, so are there more solutions or are there just 2?

OpenStudy (anonymous):

oh wow that link is awesome lol

ganeshie8 (ganeshie8):

\(\large \tan x = -\sqrt{3}\) gives you two more solutions

ganeshie8 (ganeshie8):

scroll down to see the "Solutions"

OpenStudy (anonymous):

yeah im seeing it now, thanks